OpenStudy (anonymous):

You are ordering a pizza and soda for the big game. A deep dish pizza costs $10.25, and a two-liter of soda is $2.00. For each touchdown that the Eagles score on Sunday, Vinny's Pizza lowers the cost by $1.50. How many touchdowns will the Eagles have to score for your bill to come to $4.75? Explain how to solve for the missing value using an equation.

OpenStudy (anonymous):

10.25 + 200 - 1.50x = 4.75 12.25 - 1.50x = 4.75 -1.50x = 4.75 - 12.25 -1.50x = - 7.50 x = -7.50/-1.50 x = 5 check... 12.25 - 1.5(5) = 4.75 12.25 - 7.50 = 4.75 4.75 = 4.75 (correct) 5 touchdowns is needed

OpenStudy (anonymous):

any questions ?